# Fine-grained Spatial-temporal MLP Architecture for Metro Origin-Destination Prediction

### v1

@@ 4.4. Bidirectional Trend Learner (BTL)
(9) $\boldsymbol{H_{f}}=\mathrm{Conv1d}([\boldsymbol{H_{\mathrm{prev}}};\boldsymbol{H_{\mathrm{cur}}}])$
@@
(10) $\boldsymbol{H_{g}}=\delta(\boldsymbol{W^{g}}\boldsymbol{H_{f}})$
@@
(11) $\boldsymbol{\tilde{H}_{\mathrm{prev}}}=\boldsymbol{H_{g}}\ast(\boldsymbol{W^{p}}\boldsymbol{H_{\mathrm{prev}}})+\boldsymbol{H_{\mathrm{prev}}}$
@@
(12) $\boldsymbol{\tilde{H}_{\mathrm{prev}}},\boldsymbol{\tilde{H}_{\mathrm{cur}}}=\mathrm{BTL}(\boldsymbol{H_{\mathrm{prev}}},\boldsymbol{H_{\mathrm{cur}}})$
@@
4.5. Optimization
In this subsection, we first introduce double-branch architecture of the model and then proceed to detail the model’s loss function. Subsequently, we perform time complexity analysis for each component and the model.
4.5.1. Loss Function
@@
(13) $\mathrm{Loss}$ $=L_{d-1}+L_{d}$
$L_{\{d-1,d\}}$ $=\frac{\sum_{i=1}^{N}\sum_{j=1}^{N}|\widetilde{OD}_{t+1}^{\{d-1,d\}}-OD_{t+1}^{\{d-1,d\}}|}{N^{2}}$
where $\widetilde{OD}_{t+1}^{\{d-1,d\}}$ and $OD_{t+1}^{\{d-1,d\}}$ represent the predicted and true values of the OD matrix at the $(t+1)^{th}$ time interval of $(d-1)^{th}$ or $d^{th}$ day, respectively.
@@ 4.5.2. Model Complexity
Combining these analyses, the ODMixer model demonstrates a time complexity of $O(N^{2})$. The training and inference time of both the baseline and ODMixer models are evaluated through experiments, as showed in Table 7 and 8.
5. Experiment
@@
5.1. Experimental Configuration
5.1.1. Dataset Description
We conduct experiments on two real-world datasets (Liu et al., 2022c), as shown in Table 2. HZMOD records transaction data for 80 stations from January 1, 2019, to January 25, 2019. SHMOD records transaction data for 288 stations from July 1, 2016, to September 30, 2016. These transaction data are segmented into 15-minute interval, and the complete dataset is partitioned into training, validation, and test set following a specific ratio.
Table 2. Details of the datasets.
@@ 5.1.1. Dataset Description
Table 3. Performance of OD prediction on HZMOD dataset.
@@ 5.1.2. Implementation Details
5.1.3. Quantitative Evaluation Metrics
@@
(14) $\mathrm{MAE}$ $=\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}|\widetilde{OD}(i,j)-OD(i,j)|$
@@
5.2. Comparison with State-of-the-Art Methods
We conduct a comparative analysis between our proposed model and a selection of classical as well as contemporary models, following as: Historical Average (HA), LSTM (Memory, 2010), GRU (Cho et al., 2014), Graph WaveNet (Wu et al., 2019), DCRNN (Li et al., 2018), STG2Seq (Bai et al., 2019), PVCGN (Liu et al., 2020), DGSL (Shang et al., 2021), Informer (Zhou et al., 2021), HIAM (Liu et al., 2022c). STCNN (Zhang et al., 2023b) and ST-VGCN(Yang et al., 2022) that consider OD pairs are included because they have no publicly accessible codes and have not been evaluated on identical datasets.
Table 4. Performance of OD prediction on SHMOD dataset.
@@ 5.2. Comparison with State-of-the-Art Methods
Table 5. Performance of different variants on HZMOD dataset.
@@
(b) Impact of Number of Channel
Figure 3. Impact of Hyper-parameters.
We have conducted experiments on two datasets, and the experimental results for all methods are presented in Table 3 and 4. Firstly, the simplest method, HA, can partially reflect the flow range of the training set by directly averaging it. However, lacking generalization ability, it fails to capture temporal variations in traffic. Consequently, its performance tends to be inferior to that of ODMixer, with the performance gap widening as the dataset size increases. After analysis, it is evident that the performance of LSTM, GRU, Graph WaveNet, and STG2Seq models is inferior to that of HA, potentially due to their oversimplified design that neglects the spatial-temporal dependencies among metro stations. The incorporation of spatial-temporal relations in models like DCRNN and Informer has resulted in better performance. DGSL outperforms other baselines on the SHMOD dataset because it learns the graph structure rather than using the provided stricture, which allows for a more effective capture of relations on complex dataset. PVCGN, which considers three types of graphs and comprehensively accounts for the interstation relations, exhibits good performance across both datasets. HIAM, which integrates OD and DO information and considers the incompleteness of OD matrices, achieves commendable results but performs less effectively than DGSL on the SHMOD dataset, possibly because the large scale of SHMOD requires a more fine-grained approach than can be provided by neighborhood graph aggregation alone to capture the complex interactions among numerous stations.
Table 6. Performance of different variants on SHMOD Dataset.
@@ 5.3. Ablation Study
The results are presented in the Table 5 and 6. ODMixer-w/o CM exhibits inferior performance compared to ODMixer, signifying the effectiveness of Channel Mixer in modeling short-term temporal relations of OD pairs. ODMixer-w/o MM performs less effectively than both ODMixer-w/o OM and ODMixer-w/o DM, highlighting the necessity of both Origin and Destination perspectives. The results indicate that these two perspectives complement each other in considering the relations among OD pairs. ODMixer-w/o BTL and ODMixer-w/o PB show some performance losses, suggesting that accounting for the temporal similarity of metro flows can enhance the model’s performance. Due to the difficulty of accurately modeling long-term dependencies for large-scale metro data, ODMixer performs slightly worse than ODMixer-w/o BTL in RMSE metrics. These experimental findings affirm that all modules contribute to the ODMixer.
Table 7. Efficiency study of ODMixer on HZMOD dataset.
@@
*” indicates the Batch Size in PVCGN is 16.
5.4. Hyper-parameters Analysis
@@
5.5. Efficiency Analysis
Table 8. Efficiency study of ODMixer on SHMOD dataset
@@ 5.5. Efficiency Analysis
Table 7 and 8 present a comparative analysis of all methods. The LSTM, GRU, Graph WaveNet, and STG2Seq models exhibit poorer performance than HA, despite their shorter training and inference times. Despite Informer’s short training and inference times on the HZMOD dataset, it has a substantial number of parameters. Regarding the SHMOD dataset, Informer cannot be trained , even when the batch size is 1. Furthermore, the PVCGN model is unable to be trained on a single GPU with a batch size of 32 for HZMOD, necessitating a reduction to 16 for testing. Despite its good performance, the PVCGN model requires significantly longer training and inference times. In contrast, the HIAM model, which serves as the best-performance baseline, has a higher number of parameters, as well as longer training and inference times compared to ODMixer. Our proposed ODMixer not only achieves superior performance but also maintains reasonable training and inference times.
6. Conculsion
@@
In this paper, we introduce ODMixer, a fine-grained spatial-temporal MLP architecture to address the metro OD prediction problem. Our ODMixer learns the short-term temproal relations of OD pairs by incorporating the Channel Mixer. The Multi-view Mixer efficiently captures OD pair relations from both origin and destination perspectives. With the integration of BTL, our ODMixer can perceive long-term traffic changes. Experimental results represent ODMixer’s outstanding performance on two large-scale datasets.
Future directions for ODMixer involve incorporating additional city information, such as urban population distribution, regional composition, and Point of Interest (POI), to further enrich feature representation. Moreover, enhancing the model’s ability to learn the general pattern of traffic flow can improve its migration capability and scalability. Deploying ODMixer in actual metro systems can facilitate the management and optimization of metro operations, thereby enhancing transportation efficiency.

### v2

@@ IV-D Bidirectional Trend Learner (BTL)
$\boldsymbol{H_{f}}=\mathrm{Conv1d}([\boldsymbol{H_{\mathrm{prev}}};\boldsymbol{H_{\mathrm{cur}}}])$ (9)
@@
$\boldsymbol{H_{g}}=\delta(\boldsymbol{W^{g}}\boldsymbol{H_{f}})$ (10)
@@
$\boldsymbol{\tilde{H}_{\mathrm{prev}}}=\boldsymbol{H_{g}}\ast(\boldsymbol{W^{p}}\boldsymbol{H_{\mathrm{prev}}})+\boldsymbol{H_{\mathrm{prev}}}$ (11)
@@
$\boldsymbol{\tilde{H}_{\mathrm{prev}}},\boldsymbol{\tilde{H}_{\mathrm{cur}}}=\mathrm{BTL}(\boldsymbol{H_{\mathrm{prev}}},\boldsymbol{H_{\mathrm{cur}}})$ (12)
@@
IV-E Optimization
In this subsection, we first introduce double-branch architeture of the model and then proceed to detail the model’s loss function. Subsequently, we perform time complexity analysis for each component and the model.
IV-E1 Loss Function
@@
$\mathrm{Loss}$ $=L_{prev}+L_{cur}$ (13)
$L_{prev}$ $=\frac{\sum_{i=1}^{N}\sum_{j=1}^{N}|\widetilde{OD}_{t+1}^{d-1}-OD_{t+1}^{d-1}|}{N^{2}}$
$L_{cur}$ $=\frac{\sum_{i=1}^{N}\sum_{j=1}^{N}|\widetilde{OD}_{t+1}^{d}-OD_{t+1}^{d}|}{N^{2}}$
where $\widetilde{OD}_{t+1}^{d-1}$, $\widetilde{OD}_{t+1}^{d}$ and $OD_{t+1}^{d-1}$, $OD_{t+1}^{d}$ represent the predicted and true values of the OD matrix at the $(t+1)^{th}$ time interval of $(d-1)^{th}$ or $d^{th}$ day, respectively.
@@ IV-E2 Model Complexity
Combining these analyses, the ODMixer model demonstrates a time complexity of $O(N^{2})$. The training and inference time of both the baseline and ODMixer models are evaluated through experiments, as showed in Table VII and VIII.
V Experiment
@@
V-A Experimental Configuration
V-A1 Dataset Description
We conduct experiments on two real-world datasets [5], as shown in Table II. HZMOD records transaction data for 80 stations from January 1, 2019, to January 25, 2019. SHMOD records transaction data for 288 stations from July 1, 2016, to September 30, 2016. These transaction data are segmented into 15-minute interval, and the complete dataset is partitioned into training, validation, and test set following a specific ratio.
TABLE II: Details of the datasets.
@@ V-A1 Dataset Description
TABLE III: Performance of OD prediction on HZMOD dataset.
@@ V-A2 Implementation Details
V-A3 Quantitative Evaluation Metrics
@@
$\mathrm{MAE}$ $=\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}|\widetilde{OD}(i,j)-OD(i,j)|$ (14)
@@
V-B Comparison with State-of-the-Art Methods
We conduct a comparative analysis between our proposed model and a selection of classical as well as contemporary models. STCNN [9] and ST-VGCN[10] that consider OD pairs are included because they have no publicly accessible codes and have not been evaluated on identical datasets.
  • Historical Average (HA): Future projections were derived by directly utilizing historical flow averages for each metro station.
  • LSTM[43]: This model utilizes a Seq2Seq approach, employing two LSTM layers to predict future metro OD.
  • GRU[44]: The architecture of this model closely resembles the previous one, with a key distinction being the use of GRU layers instead of LSTM layers.
  • Graph WaveNet[45]: The model acquires adaptive dependency matrices within the data by embedding nodes, and it captures long-range temporal dependencies through the use of stacked dilated 1D convolutions.
  • DCRNN[46]: DCRNN captures spatial dependencies through the utilization of bi-directional random walks on the graph. Additionally, it employs encoder-decoder structures with scheduled sampling to model temporal information.
  • STG2Seq[47]: STG2Seq incorporates a hierarchical graph convolution structure to effectively capture both spatial and temporal correlations.
  • PVCGN[48]: PVCGN comprehensively learns complex spatio-temporal relations between metro stations by constructing physical, similarity, and correlation graphs.
  • DGSL[49]: DGSL optimizes the average performance of the graph distribution by leveraging probabilistic graph models when the graph structure is unknown, thus facilitating the learning of the graph structure.
  • Informer[50]: Informer is a transformer-based model specifically designed for the task of long sequence time-series forecasting. It is notable for its efficient ProbSparse self-attention mechanism, self-attention distillation, and generative-style decoder. For our metro OD prediction, we built our model using the official code of Informer as a foundation, and we customized it to suit our specific requirements.
  • HIAM[5]: HIAM integrates diverse information to jointly learn the evolutionary patterns of OD and OD, featuring two branch for OD and DO estimation separately and a DIT for enhanced OD-DO correlation modeling.
TABLE IV: Performance of OD prediction on SHMOD dataset.
@@ V-B Comparison with State-of-the-Art Methods
TABLE V: Performance of different variants on HZMOD dataset.
@@
(b) Impact of Number of Dimension
Figure 4: Impact of Hyper-parameters. The two figure illustrate the impact of different parameters on prediction results across two datasets. Figure (a) displays the effect of varying the number of layers on model performance, while Figure (b) shows the influence of different feature dimensions on model performance.
We have conducted experiments on two datasets, and the experimental results for all methods are presented in Table III and IV. Firstly, the simplest method, HA, can partially reflect the flow range of the training set by directly averaging it. However, lacking generalization ability, it fails to capture temporal variations in traffic. Consequently, its performance tends to be inferior to that of ODMixer, with the performance gap widening as the dataset size increases. After analysis, it is evident that the performance of LSTM, GRU, Graph WaveNet, and STG2Seq models is inferior to that of HA, potentially due to their oversimplified design that neglects the spatial-temporal dependencies among metro stations. The incorporation of spatial-temporal relations in models like DCRNN and Informer has resulted in better performance. DGSL outperforms other baselines on the SHMOD dataset because it learns the graph structure rather than using the provided stricture, which allows for a more effective capture of relations on complex dataset. PVCGN, which considers three types of graphs and comprehensively accounts for the interstation relations, exhibits good performance across both datasets. HIAM, which integrates OD and DO information and considers the incompleteness of OD matrices, achieves commendable results but performs less effectively than DGSL on the SHMOD dataset, possibly because the large scale of SHMOD requires a more fine-grained approach than can be provided by neighborhood graph aggregation alone to capture the complex interactions among numerous stations.
TABLE VI: Performance of different variants on SHMOD Dataset.
@@ V-C Ablation Study
The results are presented in the Table V and VI. ODMixer-w/o CM exhibits inferior performance compared to ODMixer, signifying the effectiveness of Channel Mixer in modeling short-term temporal relations of OD pairs. ODMixer-w/o MM performs less effectively than both ODMixer-w/o OM and ODMixer-w/o DM, highlighting the necessity of both Origin and Destination perspectives. The results indicate that these two perspectives complement each other in considering the relations among OD pairs. ODMixer-w/o BTL and ODMixer-w/o PB show some performance losses, suggesting that accounting for the temporal similarity of metro flows can enhance the model’s performance. Due to the difficulty of accurately modeling long-term dependencies for large-scale metro data, ODMixer performs slightly worse than ODMixer-w/o BTL in RMSE metrics. These experimental findings affirm that all modules contribute to the ODMixer.
TABLE VII: Efficiency study of ODMixer on HZMOD dataset.
@@
    The Batch Size in PVCGN is 16.
V-D Hyper-parameters Analysis
@@
V-E Efficiency Analysis
TABLE VIII: Efficiency study of ODMixer on SHMOD dataset
@@ V-E Efficiency Analysis
Table VII and VIII present a comparative analysis of all methods. The LSTM, GRU, Graph WaveNet, and STG2Seq models exhibit poorer performance than HA, despite their shorter training and inference times. Despite Informer’s short training and inference times on the HZMOD dataset, it has a substantial number of parameters. Regarding the SHMOD dataset, Informer cannot be trained , even when the batch size is 1. Furthermore, the PVCGN model is unable to be trained on a single GPU with a batch size of 32 for HZMOD, necessitating a reduction to 16 for testing. Despite its good performance, the PVCGN model requires significantly longer training and inference times. In contrast, the HIAM model, which serves as the best-performance baseline, has a higher number of parameters, as well as longer training and inference times compared to ODMixer. Our proposed ODMixer not only achieves superior performance but also maintains reasonable training and inference times.
V-F Validity Verification
Since our method is based on the perspective of OD pairs, while most existing methods are station-based, we aim to verify whether our model structure can effectively address the issues in existing methods from the same perspective. To this end, we replaced the ODIM module in ODMixer with a standard Attention module, naming the modified model ODTrans. We then conducted experiments, and the results are presented in the following Table IX.
TABLE IX: Performance of OD prediction on HZMOD dataset.
Models wMAPE ($\downarrow$) Train (s) Test (s) Layer Dim BatchSize GPUs
$\text{ODTrans}^{1}$ 47.422% 19.02 3.92 2 16 8 8
ODMixer 40.609% 6.75 2.15 2 16 32 1
$\Delta$ 14,37% ×$\times 2.82$$\times 1.82$
$\text{ODTrans}^{1}$ 47.007% 27.80 4.70 3 16 8 8
ODMixer 40.781% 8.04 2.59 3 16 32 1
$\Delta$ 13.24% ×$\times 3.46$$\times$ 1.81 - - - -
  • 1
    For the Transformer, we set the batch size to 8, the number of layers to 2 or 3, and use 8 GPUs to ensure proper execution.
As described in the introduction, the number of input tokens is $N^{2}$ from the perspective of OD pairs, resulting in space and time complexities of $N^{4}$ for the attention mechanism. Consequently, ODTrans cannot run with the same number of GPUs and batch size as ODMixer when using the same model layers and dimensions. For a fair performance comparison, we reduced the number of model layers and batch size and employed multiple GPUs for ODTrans. We also used the same number of layers for training ODMixer to ensure fairness.
The experiments show that the performance of ODMixer is improved by 14.37% and 13.24% at layers 2 and 3 compared to ODTrans, respectively. Additionally, by comparing training and testing times, we found that ODMixer’s training time and testing time at layer 2 were accelerated by 2.82 times and 1.82 times, respectively. At layer 3, training and testing times were accelerated by 3.46 times and 1.81 times, respectively. These results demonstrate that the ODMixer module, designed specifically for the metro OD prediction task, effectively captures the spatiotemporal relations in traffic. This significantly enhances the model’s operational efficiency, reduces the required space, and greatly improves its deployability.
V-G Case Study
Figure 5: Case for SHMOD Dataset. It shows the prediction results of the DGSL and ODMixer models from September 10 to 13, with the 12th being a Monday. It can be observed that when traffic flow changes rapidly, the performance of DGSL is significantly worse than that of ODMixer.
In this subsection, we use a practical example to demonstrate why our ODMixer achieves better results.
As shown in Fig. 5, the OD traffic on Saturday and Sunday is minimal and stable, allowing both DGSL and ODMixer to fit the real traffic well. However, during the peak periods of Monday and Tuesday, the traffic between OD pairs changes rapidly. Although DGSL senses the growth and attenuation of traffic, its estimated peak value remains small. Conversely, ODMixer quickly detects these rapid changes and accurately estimates the peak value of the OD pair’s traffic. This is mainly due to ODMixer’s input modeling based on the OD pair perspective. The Channel Mixer captures the short-term temporal relations, the Multi-view Mixer learns the spatial relations between OD pairs, and BTL senses long-term traffic similarity. These designs enable ODMixer to handle rapid changes and complex relations in traffic effectively, significantly improving the model’s prediction accuracy and adaptability.
VI Conclusion
@@
In this paper, we introduce ODMixer, a fine-grained spatial-temporal MLP architecture to address the metro OD prediction problem. Our ODMixer learns the short-term temporal relations of OD pairs by incorporating the Channel Mixer. The Multi-view Mixer efficiently captures OD pair relations from both origin and destination perspectives. With the integration of BTL, our ODMixer can perceive long-term traffic changes. Experimental results represent ODMixer’s outstanding performance on two large-scale datasets. Future directions for ODMixer involve incorporating additional city information, such as urban population distribution, regional composition, and Point of Interest (POI). Moreover, enhancing the model’s ability to learn the general pattern of traffic flow can improve its migration capability and scalability.